OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4

OpenStudy (helder_edwin):

you have typos in the question u posted

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

i do?

OpenStudy (helder_edwin):

what you wrote has no equations!

OpenStudy (anonymous):

this is the question my teacher gavve me Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4

OpenStudy (helder_edwin):

first. are your equations \[ \large y=x-5 \] \[ \large y=-x-4 \]

OpenStudy (helder_edwin):

which points are u given?

OpenStudy (anonymous):

(-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (helder_edwin):

ok you have to replace the points and see if the equations are satisfied do u know how?

OpenStudy (anonymous):

No :/ im taking an online math course and i feel stpid cause i dont get any of this

OpenStudy (helder_edwin):

don't say that. if you keep saying that one day you'll believe it!

OpenStudy (helder_edwin):

give a moment

OpenStudy (anonymous):

Thanks..Im about to take a math quiz do you wanna help me through it?

OpenStudy (helder_edwin):

were i am it is late

OpenStudy (helder_edwin):

let's do this one ok?

OpenStudy (helder_edwin):

i'll draw it

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):